A potential difference of 2.00 V will be applied to a 39.00 m length of 18-gauge aluminum wire (diameter = 0.0400inches). Calculate the current. Calculate the current density. Calculate the electric field. Calculate the rate at which thermal energy will appear in the wire.
Answer #1

V = iR

i = V/R
i = 3.8/R

to calculate R ,
Resistance = resistivity x length/area
R = ?L/A =0.223

i = 3.8 / 0.223 = 17 A